OpenStudy (anonymous):

if the sum and a number and two is tripled, the result is nine less than twice the number. find the number

OpenStudy (radar):

Not for sure I am following this question. Let N = the Number N + 2 = sum of the number and 2 3(N+2) us three times the sum 2N = twice the number 2N - 9 is twice the number - 9 3(N + 2) = 2N - 9 Take it from there.